# Is animal magnetoreception quantum? A perspective from the energy resolution limit

### v1

@@ III.2 What is $\delta B$ in vivo?
The aforementioned numbers involve numerous observations with multiple species, therefore it is virtually impossible to add an error bar weighting the above estimates. Hence we here use the average of all of the above estimates and set $\delta B=0.3~{}{\rm\mu T}$.
@@ III.3 Energy resolution in vivo: cryptochrome
A prevalent model [138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151] for radical-pair based magnetoreception has been the radical-pair reaction following electron transfer from a triad of tryptophan residues (Tr) to flavin adenine dinucleotide (FAD), with the radical-pair ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ embedded in the protein cryptochrome (Cry), in particular Cry4. This protein is expressed in the retina of the European robin and exhibits strong binding with flavin [6, 152, 153, 154]. Recently a tetrad of tryptophans was considered [147], and we will here use this model, depicted in Fig. 3. In the same figure we depict the known distances between the radicals, from which we can estimate the volume of the sensor as $V=2.1~{}{\rm nm}\times(0.8~{}{\rm nm})^{2}=1.4\times 10^{-27}~{}{\rm m}^{3}$.
Figure 3: Cryptochrome protein (shaded grey background) is the host of the ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ radical-pair working as a magnetic sensor. Center-to-center distances between FAD and tryptophans are known and provide the overall volume of the sensor. We estimate the length of the sensor from the FAD-W369 distance ($21.3~{}{\rm\AA}$) and the width from the W372-W369 distance projected on the $y$-axis ($8~{}{\rm\AA}$). Figure reproduced from [147] under the Creative Commons Attribution License.
@@
Lastly, we need to estimate the measurement time $\tau$, given by the lifetime of the ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ radical-pair. The lifetime of ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ in vivo is not yet known. In vitro measurements at $1~{}{\rm{}^{\circ}C}$ report $6~{}{\rm\upmu s}$ [155, 156]. As suggested in [157], a lower bound on $\tau$ is given by the electron spin precession time in earth’s field, which is $0.7~{}\upmu{\rm s}$. The authors in [158] find an operational window of the compass for $\tau=10~{}\upmu{\rm s}$, given the relaxation produced by rotational motion of the radical-pair, in particular FAD, and taking into account the sharp “needle” described in [46]. The latter is an impressively sharp feature appearing in the angular dependence of the reaction yields for lifetimes larger than $\tau=5~{}\upmu{\rm s}$, when the realistic hyperfine tensors for ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ are used. If we allow ourselves to hypothesize, it seems unlikely that Nature did not evolve to take advantage of this feature. Based on all of the above, we here set $\tau=10~{}\upmu{\rm s}$.
@@
Putting it altogether, it follows that ${\rm ER}=4.8\times 10^{-6}\hbar$. Is the ERL violated? Not at all. The previous result shows that the signal from at least $N_{\rm cry}^{0}=1/(4.8\times 10^{-6})\approx 2\times 10^{5}$ cryptochrome proteins must be physiologically integrated in order to obtain the magnetic sensitivity $\delta B$ within the time $\tau$. This is exactly the utility of the ERL, it informs the workings of this magnetoreception mechanism in model-independent ways. If, for example, it is found that indeed there are only $N_{\rm cry}^{0}$ proteins contributing, then cryptochrome magnetoreception will be working right at the quantum limit. On the other hand, it is found that there are e.g. ten times as many proteins contributing, then cryptocrhome magnetoreception will be working at the level of $10\hbar$, and there is in principle room for a ten-fold improvement, were one to mimic Nature.
@@
Overall, the values for $\widetilde{\delta B}=\delta B\sqrt{\tau}$ and the spatial dimension $L$ that we will use for cryptochrome magnetoreception in our global results of Sec. VII are $\widetilde{\delta B}\approx 1~{}{\rm nT/\sqrt{Hz}}$ and $L=(N_{\rm cry}V)^{1/3}\geq 66~{}{\rm nm}$, where $N_{\rm cry}$ the number of cryptochrome proteins contributing to the estimate of the magnetic field, and necessarily, it is $N_{\rm cry}\geq N_{\rm cry}^{0}$.
@@ III.4 Comment on the sensor volume
One might rightly wonder why in the discussion of CPF in Sec. III.A we considered the whole volume of the solution probed by the laser, whereas in the discussion of cryptochrome in Sec. III.C we counted individual sensors, i.e. we multiplied the volume of the ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ sensor with the number of such sensors, which is the number of cryptochrome proteins involved in magnetoreception. An analogy with optical pumping magnetometers [129] will facilitate this discussion.
@@
Such magnetometers, schematically shown in Fig. 4a, involve atomic spins in the vapor phase, with the atomic vapor enclosed in a glass cell. The atoms’ spin state, conveying information about the magnetic field, is probed by a laser. One can use a photodiode detecting the whole laser beam, alternatively, one can use a photodiode array (Fig. 4b). The array consists of $n$ photosensitive elements. If one detects the individual signals $s_{1}$, $s_{2}$, …, $s_{n}$, then each signal conveys information about the magnetic field experienced by the atoms inside the volume defined by the area of one element and the laser path length through the vapor (Fig. 4c). If, on the other hand, the $n$ individual signals are integrated (Fig. 4d), information about the magnetic field within the aforementioned elementary volumes is lost, and the sensitive volume is now defined by the laser path length in the vapor and the whole area of the photodiode array (essentially, this is the total vapor volume, when the laser beam waist covers the whole cell).
@@
Figure 4: (a) Sensitive volume in atomic vapor magnetometers. An ensemble of spins in an atomic vapor is enclosed in a cell and probed by a laser, which is detected by (b) a photodiode array. (c) If one measures the individual signals, $s_{1}$, $s_{2}$, …, $s_{n}$, produced by the array, then the sensitive volume is determined by the area of one photosensitive element and the laser path length in the vapor. (d) If one integrates the $n$ signals produced by the array, the sensitive volume is determined by the whole area of the photodiode array and the path length, essentially the total volume of the vapor (if the laser waist covers the whole cell). (e) In the retina, rod cells are similarly integrated by bipolar cells (figure reproduced from [159] under the Creative Commons Attribution License).
Biological systems have devised such integration mechanisms. For example (Fig. 4e), rod cells in the retina are integrated by bipolar cells. Going back to magnetoreception, if a number of magnetoreceptive elements are integrated and thus contribute as a whole to the biochemical signal conveying a magnetic field change ($B\rightarrow B+\delta B$) to the animal’s brain, then the sensitive volume is the volume of each element times their number. But then, in atomic sensors why isn’t the sensitive volume defined by the atomic volume (about $1~{}{\rm\AA}^{3}$) times the number of atoms in the vapor, but instead [85], it is defined by the volume occupied by the whole vapor ?
A more fundamental way to address this issue is the following. In the derivation of the ERL presented in [104], it is shown that it is the information flow between “system” and “meter” that plays a crucial role. In quantum measurements, two parties are involved in the measurement process, the “system” e.g. the spin, and another physical object interacting with the system, called the “meter”. Due to their interaction, information about the system’s state is encoded in the meter. The information flow between system and meter takes place inside the sensitive volume during the measurement time. In atomic vapors the meter is another atom of the vapor itself, with every atom experiencing many collisions with other atoms and exchanging information about each others’ spin state, while relaxing the spin. It is thus all kinds of binary collisions within the whole vapor that drive these effects. The probing laser cannot resolve these local effects, hence the sensitive volume is defined by the volume of the whole vapor.
Equivalently, within the spin coherence time one atom in the vapor samples a volume much larger than the atomic volume itself. For example, such vapors usually include a buffer gas to eliminate spin relaxation at the cell walls. Even for buffer gas pressures as large as e.g. 5 atm, the diffusion constant of alkali vapors in e.g. nitrogen is $D\approx 0.06~{}{\rm cm^{2}/s}$, thus the atom samples a volume of $(DT_{2})^{3/2}\approx 10^{-22}~{}{\rm m^{3}}$ within the spin coherence time of about $T_{2}\approx 10~{}{\rm ms}$. This volume is 8 orders of magnitude larger than the atomic volume.
In the case of radical-pairs in cryptocrhome, it is the reaction products that convey the information on the magnetic field, and these are initially localized in the volume defined earlier. The ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ radical-pair is also localized in its protein host. The reaction products could diffuse elsewhere, and they might trigger many more information transfer processes, but fundamentally, the termination of the measurement is the end of the reaction, taking place within the volume defined by ${\rm FAD^{\bullet-}Tr^{\bullet+}}$. Put differently, it is hard to suggest any “non-local” interactions that would affect spin dynamics of one ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ radical-pair hosted in one cryptochrome, with another pair hosted in another cryptochrome, which is about 10 nm away (the size of cryptochrome, 5 times as large as the size of ${\rm FAD^{\bullet-}Tr^{\bullet+}}$). This is because interactions like exchange or dipolar fade away with distance fast enough [160] to have an appreciable inter-cryptochrome effect.
Even if we consider a free radical at the “outskirts” of cryptochrome, i.e. a distance of 10 nm away from the ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ pair, the dipolar magnetic field produced would be about $10~{}\upmu{\rm T}$, comparable to earth’s field, hard to reconcile with a working magnetoreceptor. In a nutshell, the assumption that spin relaxation in ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ is local seems rather likely. In such case, the magnetoreceptor’s volume is $N_{\rm cry}$ times the volume of the radical-pair molecule. Essentially, the total volume scaling as $N_{\rm cry}$ means that the magnetic sensitivity $\delta B$ scales as $1/\sqrt{N_{\rm cry}}$, which is the standard statistical gain over many independent repetitions of a measurement.
@@
In the case of CPF, on the other hand, the typical distance between molecules is 25 nm, while the size of CPF is about 8 nm, and solvation effects in electron-transfer dynamics can make non-local effects more pronounced compared to cryptocrhome. The probing laser does not resolve such effects, thus we defined the sensitive volume with the volume of the solution probed by the laser.
@@ IV Magnetite Mechanism
Figure 5: (a) Axon bundle with several iron containing dendrites (figure reproduced from [176] under the Creative Commons Attribution License). (b) Magnetoreception model based on a chain of superparamagnetic clusters attached to a membrane [183]. (c) When the magnetic field is non-zero and points along the chain, the clusters get magnetized along the same direction and attract each other, deforming the membrane and conveying the magnetic field change by changing ion-channel currents. If the magnetic field points perpendicular to the chain (not shown here), the clusters repel each other. (d) Geometry of the MagR-cryptochrome protein complex as derived from the measurements presented in [56].
@@
To evaluate the measurement time $\tau$ we neglect the elastic properties of the filaments attaching the clusters to the membrane and consider just two forces acting on the spherical cluster, the magnetic force $F_{m}$ and the drag force of the cytoplasm $F_{d}$. The latter is given by $F_{d}=6\pi\eta_{c}rv$, where $\eta_{c}\approx 3.5\times 10^{-3}~{}{\rm Ns/m^{2}}$ is the cytoplasm viscosity, known to be 5 times that of water [186], $v$ the sphere’s velocity, and $r$ its radius. The sphere, of mass density about $\rho_{s}=5~{}{\rm g/cm^{3}}$, reaches its terminal velocity, $v=F_{m}/6\pi\eta_{c}r\approx 3~{}{\rm\upmu m/s}$, within time $2\rho_{s}r^{2}/9\eta_{c}\approx 30~{}{\rm ns}$.
@@ V MagR
The authors in [56, 191, 192, 193, 194] used genome-wide screening and a tour-de-force experimental validation to discover an iron-sulphur cluster protein, called MagR, which forms a rod-like complex with cryptochromes. The geometry of the rod is shown in Fig. 5d. It has a length of 25 nm and diameter of 12 nm, thus volume $V=3\times 10^{-24}~{}{\rm m^{3}}$. The workings of this complex are not yet understood. It might turn out that the cryptochromes work synergistically with MagR, in some sort of synthesis of the two mechanisms considered previously. We here make this working assumption. Since (i) the biophysics of the MagR complex remains to be unravelled, and (ii) the cryptochrome magnetoreception has a $\tau$ about one order of magnitude smaller than the $\tau$ of the magnetite mechanism, we here use as measurement time $\tau$ the geometric mean of these two numbers, i.e. $\tau=80~{}\upmu{\rm s}$. The resulting ER is $0.08\hbar$, meaning that just $N_{\rm MagR}^{0}=1/0.08=12$ complexes need to be integrated for the mechanism to hit the ERL bound.
@@
Figure 6: Magnetic sensitivity $\widetilde{\delta B}=\delta B\sqrt{\tau}$ as a function of $L$, the linear dimension of the sensor, for several mechanisms considered in animal magnetoreception. The red solid line defines the energy resolution limit given by $\widetilde{\delta B}=\sqrt{2\mu_{0}\hbar/L^{3}}$. Cryptochrome magnetoreception based on the FAD-Tr radical-pair sits right at the limit when $N_{\rm cry}^{0}=2\times 10^{5}$ cryptochromes contribute to a magnetic sensitivity of $\delta B=0.3~{}\upmu{\rm T}$ within time $\tau=10~{}\upmu{\rm s}$. If the number of contributing cryptochromes is e.g. $10^{3}$ times larger than $N_{\rm cry}^{0}$, the total volume of the radical-pair sensor, and thus the ERL, also increases by $10^{3}$, while the linear dimension by 10. This is depicted by the green/blue arrows pointing towards increasing linear dimension. The blue arrow shows the ER in units of $\hbar$, and the green arrow the ratio $N_{\rm cry}/N_{\rm cry}^{0}$. Similar comments apply to the MagR sensor, shown to hit the ERL when only $N_{\rm MagR}^{0}=12$ MagR-protein complexes contribute to detecting the same $\delta B$ within $\tau=80~{}\upmu{\rm s}$. For the magnetite mechanism involving a chain of superparamagnetic clusters it is $\widetilde{\delta B}=5~{}{\rm nT/\sqrt{Hz}}$ for a linear dimension of $3~{}\upmu{\rm m}$. The laser spectroscopic measurement of CPF takes place in a solution of 3 mm linear dimension, with the corresponding sensitivity being $\widetilde{\delta B}=2.5~{}{\rm nT/\sqrt{Hz}}$. Picture of cryptochrome protein adapted from [149] under the Creative Commons Attribution License.
@@
Overall, the values for $\widetilde{\delta B}=\delta B\sqrt{\tau}$ and the spatial dimension $L$ that we will use for MagR magnetoreception in our global results of Sec. VII are $\widetilde{\delta B}\approx 2~{}{\rm nT/\sqrt{Hz}}$ and $L=(N_{\rm MagR}V)^{1/3}\geq 33~{}{\rm nm}$, where $N_{\rm MagR}$ is the number of MagR complexes contributing to the estimate of the magnetic field, and necessarily, it is $N_{\rm MagR}\geq N_{\rm MagR}^{0}$.

### v2

@@ III.2 What is $\delta B$ in vivo?
The aforementioned numbers involve numerous observations with multiple species, therefore it is virtually impossible to add an error bar weighting the above estimates. Hence we here average all of the above estimates and set $\delta B=0.3~{}{\rm\mu T}$.
@@ III.3 Energy resolution in vivo: cryptochrome
A prevalent model [138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151] for radical-pair based magnetoreception has been the radical-pair reaction following electron transfer from a triad of tryptophan residues (Tr) to flavin adenine dinucleotide (FAD), with the radical-pair ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ embedded in the protein cryptochrome (Cry), in particular Cry4. This protein is expressed in the retina of the European robin and exhibits strong binding with flavin [6, 152, 153, 154]. Recently a tetrad of tryptophans was considered [147], and we will here use this model, depicted in Fig. 3. In the same figure we depict the known distances between the radicals, from which we can estimate the volume of the sensor as $V_{\rm rp}=2.1~{}{\rm nm}\times(0.8~{}{\rm nm})^{2}=1.4\times 10^{-27}~{}{\rm m}^{3}$. However, one can argue that the volume to be used in the estimate of the ER is the Cry volume, which is [144] $V_{\rm cry}\approx 5\times 10^{-25}~{}{\rm m}^{3}$. We will here use as sensitive volume the radical-pair volume, i.e. set $V=V_{\rm rp}$, and elaborate further on this issue in Sec. III.D.
[…]
@@
Lastly, we need to estimate the measurement time $\tau$, given by the lifetime of the ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ radical-pair. The lifetime of ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ in vivo is not yet known. In vitro measurements at $1~{}{\rm{}^{\circ}C}$ report $6~{}{\rm\upmu s}$ [155, 156]. As suggested in [157], a lower bound on $\tau$ is given by the electron spin precession time in earth’s field, which is $0.7~{}\upmu{\rm s}$. The authors in [158] find an operational window of the compass for $\tau=10~{}\upmu{\rm s}$, given the relaxation produced by rotational motion of the radical-pair, in particular FAD, and taking into account the sharp “needle” described in [46]. The latter is an impressively sharp feature appearing in the angular dependence of the reaction yields for lifetimes larger than $\tau=5~{}\upmu{\rm s}$, when the realistic hyperfine tensors for ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ are used. If we allow ourselves to hypothesize, it seems unlikely that Nature did not evolve to take advantage of this feature. Based on all of the above, it appears reasonable to set $\tau=10~{}\upmu{\rm s}$. Towards a more general treatment we introduce the dimensionless factor ${\tilde{\tau}}={\tau\over{10~{}\upmu{\rm s}}}$, and let $\tau$ vary between $1~{}\upmu{\rm s}$ and $100~{}\upmu{\rm s}$, thus ${\tilde{\tau}}$ ranges from 0.1 to 10.
@@
Putting it altogether, it follows that ${\rm ER}=4.8\times 10^{-6}{\tilde{\tau}}\hbar$. Is the ERL violated? Not at all. The previous result shows that the signal from at least $N_{\rm cry}^{0}(\tilde{\tau})=1/(4.8\times 10^{-6}{\tilde{\tau}})\approx 2\times 10^{5}/\tilde{\tau}$ Cry proteins must be physiologically integrated in order to obtain the magnetic sensitivity $\delta B$ within the time $10{\tilde{\tau}}~{}\upmu{\rm s}$. This is exactly the utility of the ERL, it informs the workings of this mechanism in model-independent ways, using as input a few basic parameters. If, for example, it is found that indeed there are only $N_{\rm cry}^{0}(\tilde{\tau})$ proteins contributing, then Cry magnetoreception will be working right at the quantum limit. On the other hand, if it is found that there are e.g. ten times as many proteins contributing, then Cry magnetoreception will be working at the level of $10\hbar$, and there is in principle room for a ten-fold improvement, were one to mimic Nature.
@@
Overall, the values for $\widetilde{\delta B}=\delta B\sqrt{\tau}$ and the spatial dimension $L$ that we will use for Cry magnetoreception in our global results of Sec. VII are $\widetilde{\delta B}\approx\sqrt{{\tilde{\tau}}}~{}{\rm nT/\sqrt{Hz}}$ and $L=(N_{\rm cry}V)^{1/3}\geq 65~{}{\rm nm}/\tilde{\tau}^{1/3}$, where $N_{\rm cry}$ is the number of Cry proteins contributing to the estimate of the magnetic field, and necessarily, it is $N_{\rm cry}\geq N_{\rm cry}^{0}(\tilde{\tau})$.
@@ III.4 Comment on the sensor volume
Defining the sensor volume of the radical-pair magnetoreceptor is not as straightforward as it might seem, we therefore will elaborate on this issue from a number of different perspectives.
@@
An analogy with optical pumping magnetometers [129] will facilitate this discussion. Such magnetometers, schematically shown in Fig. 4a, involve atomic spins in the vapor phase, with the atomic vapor enclosed in a glass cell. The atoms’ spin state, conveying information about the magnetic field, is probed by a laser. One can use a photodiode detecting the whole laser beam, alternatively, one can use a photodiode array (Fig. 4b). The array consists of $n$ photosensitive elements. If one detects the individual signals $s_{1}$, $s_{2}$, …, $s_{n}$, then each signal conveys information about the magnetic field experienced by the atoms inside the volume defined by the area of one element and the laser path length through the vapor (Fig. 4c). If, on the other hand, the $n$ individual signals are integrated (Fig. 4d), information about the magnetic field within the aforementioned elementary volumes is lost, and the sensitive volume is now defined by the laser path length in the vapor and the whole area of the photodiode array (essentially, this is the total vapor volume, when the laser beam waist covers the whole cell). We note that this physical picture assumes the presence of a buffer gas used to hinder atomic collisions with the walls, in particular, with a buffer gas pressure such that the diffusion time through the aforementioned elementary volumes is not shorter than the measurement time.
@@
Figure 4: (a) Sensitive volume in atomic vapor magnetometers. An ensemble of spins in an atomic vapor is enclosed in a cell and probed by a laser, which is detected by (b) a photodiode array. (c) If one measures the individual signals, $s_{1}$, $s_{2}$, …, $s_{n}$, produced by the array, then the sensitive volume is determined by the area of one photosensitive element and the laser path length in the vapor. (d) If one integrates the $n$ signals produced by the array, the sensitive volume is determined by the whole area of the photodiode array and the path length, essentially the total volume of the vapor (if the laser waist covers the whole cell).
Biological systems have devised integration mechanisms in analogy with the aforementioned photodiode array. For example, rod cells in the retina are integrated by bipolar cells, the receptive field of the latter resulting from a sort of superposition of the receptive fields of the former [159]. It is thus not surprising that radical-pair magnetoreception also involves the integrated response of a number of magnetoreceptive elements, which contribute to the biochemical signal conveying a magnetic field change ($B\rightarrow B+\delta B$) to the animal’s brain. This is necessitated by the fact that one molecule alone cannot provide the required magnetic sensitivity, which finally is obtained by a statistical averaging of the responses of $N_{\rm cry}$ radical-pairs residing in $N_{\rm cry}$ Cry proteins. In the previous subsection the volume attributed to those $N_{\rm cry}$ elementary sensors was $N_{\rm cry}V_{\rm rp}$. This choice should be accompanied by a disclaimer.
In the derivation of the ERL presented in [104], it is shown that the information flow between “system” and “meter” plays a crucial role. In quantum measurements, two parties are involved in the measurement process, the “system” e.g. the spin, and another physical object interacting with the system, called the “meter”. Due to their interaction, information about the system’s state is encoded in the meter. The information flow between system and meter takes place inside the sensitive volume during the measurement time. The measurement time is determined by spin relaxation effects. Thus the sensor volume should be the physical volume where those relaxation effects terminating the radical-pair reaction are unfolding. One can readily argue that such spin relaxation effects (related to electron transfer, molecular rotation, exchange and dipolar interactions between the two unpaired electrons of the radical-pair [160]) are localized within the radical-pair volume $V_{\rm rp}$. The reaction products could diffuse elsewhere, and they might trigger many more information transfer processes at larger spatial scales, but fundamentally, the termination of the measurement is the end of the reaction, taking place within the volume $V_{\rm rp}$ defined by ${\rm FAD^{\bullet-}Tr^{\bullet+}}$.
However, one could also imagine “non-local” interactions at the spatial scale of the whole Cry that could affect spin dynamics of the ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ radical-pair. For example, one could consider a free radical at the “outskirts” of Cry, i.e. a distance of about 10 nm away from the ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ pair. The dipolar magnetic field of the free radical would produce at the site of ${\rm FAD^{\bullet-}Tr^{\bullet+}}$ a magnetic field of about $10~{}\upmu{\rm T}$, comparable to earth’s field. This seems hard to reconcile with a working magnetoreceptor, on the other hand, the cumulative effect of several such free radicals could be suppressed, and lead to non-local relaxation consistent with the magnetic sensitivity of the sensor. Another possibility that has been touched upon in the literature is the role of hyperfine relaxation from the protein’s proton spins [161], which effect is not yet broadly understood.
Thus, the disclaimer is that while we use $V=V_{\rm rp}$, a more refined understanding of spin-relaxation in such reactions will not only constrain $\tilde{\tau}$, but might also lead to $V>V_{\rm rp}$. Either possibility will translate into a different $N_{\rm cry}^{0}$. But in any case, with $N_{\rm cry}$ radical-pairs collectively producing the magnetic sensing response, the sensor volume scales as $N_{\rm cry}$, meaning that the magnetic sensitivity $\delta B$ scales as $1/\sqrt{N_{\rm cry}}$. This is the standard statistical gain over many independent repetitions of a measurement [129], usually referred to as standard quantum limit (see also [95]).
[…]
@@
In the case of CPF in solution, the typical distance between molecules is 25 nm, while the size of CPF is about 8 nm. We here expect solvation effects in electron-transfer dynamics to make non-local spin relaxation effects non-neglibigle. The probing laser does not spatially resolve spin relaxation effects at the scale of one molecule or its solvation cage, thus we defined the sensitive volume with the total volume of the solution probed by the laser. This is equivalent to using the cage volume of $(25~{}{\rm nm})^{3}$ per molecule and multiplying by the number of molecules.
@@ IV Magnetite Mechanism
Figure 5: (a) Axon bundle with several iron containing dendrites (figure reproduced from [177] under the Creative Commons Attribution License). (b) Magnetoreception model based on a chain of superparamagnetic clusters attached to a membrane [184]. (c) When the magnetic field is non-zero and points along the chain, the clusters get magnetized along the same direction and attract each other, deforming the membrane and conveying the magnetic field change by changing ion-channel currents. If the magnetic field points perpendicular to the chain (not shown here), the clusters repel each other. (d) Geometry of the MagR-Cry protein complex as derived from the measurements presented in [56].
@@
To evaluate the measurement time $\tau$ we neglect the elastic properties of the filaments attaching the clusters to the membrane and consider just two forces acting on the spherical cluster, the magnetic force $F_{m}$ and the drag force of the cytoplasm $F_{d}$. The latter is given by $F_{d}=6\pi\eta_{c}rv$, where $\eta_{c}\approx 3.5\times 10^{-3}~{}{\rm Ns/m^{2}}$ is the cytoplasm viscosity, known to be 5 times that of water [187], $v$ the sphere’s velocity, and $r$ its radius. The sphere, of mass density about $\rho_{s}=5~{}{\rm g/cm^{3}}$, reaches its terminal velocity, $v=F_{m}/6\pi\eta_{c}r\approx 3~{}{\rm\upmu m/s}$, within time $2\rho_{s}r^{2}/9\eta_{c}\approx 80~{}{\rm ns}$.
@@ V MagR
The authors in [56, 192, 193, 194, 195, 196, 197, 198, 199] used genome-wide screening and a tour-de-force experimental validation to discover an iron-sulphur cluster protein, called MagR, which forms a rod-like complex with cryptochromes. The MagR complex has been studied further [200, 201, 202], and is already used in magnetogenetic applications [203, 204]. The geometry of the MagR complex is shown in Fig. 5d. It has a length of 25 nm and diameter of 12 nm, thus volume $V=3\times 10^{-24}~{}{\rm m^{3}}$. The exact workings of this complex are not yet understood. It seems reasonable to assume that Cry proteins work synergistically with MagR, in some sort of synthesis of the two mechanisms considered previously. We here make this working assumption.
Since the biophysics of the MagR complex remains to be unravelled, a hand-waving approach for estimating the measurement time $\tau$ would be to use the geometric mean of the values of the two previous mechanisms, Cry ($10~{}\upmu{\rm s}$), and magnetite ($300~{}\upmu{\rm s}$). Thus $\tau\approx 60~{}\upmu{\rm s}$. A more quantitative approach could follow if we use the scenario [192] that the iron-sulphur complex relates to the electron transfer processes in Cry. From the geometry in Fig. 5d we estimate the distance between the iron-sulphur complex and the Cry to be somewhere between 3 nm and 4 nm. At such long distances, we find [205, 206] electron transfer rates that range from $10^{3}~{}{\rm s}^{-1}$ to $10^{5}~{}{\rm s}^{-1}$. We thus set $\tau=100~{}\upmu{\rm s}$. Towards a more general treatment we introduce the dimensionless factor $\tilde{\tau}={\tau\over{100~{}\upmu{\rm s}}}$, and let $\tau$ vary between $10~{}\upmu{\rm s}$ and $1000~{}\upmu{\rm s}$, thus $\tilde{\tau}$ ranges from 0.1 to 10.
The resulting ER is $0.1\tilde{\tau}\hbar$, meaning that $N_{\rm MagR}^{0}(\tilde{\tau})=1/(0.1\tilde{\tau})=10/\tilde{\tau}$ complexes need to be integrated for the mechanism to hit the ERL bound. It is clear that due to the much larger volume of the complex compared to the case of radical-pair magnetoreception, fewer units are required to be right on the ERL, thus fewer Cry proteins overall. Equivalently, for a limited number of complex constituents, $\delta B$ decreases when the sensitive volume of the complex doing the sensing increases. We could imagine that this might incur some evolutionary advantage of this particular design, however, it is rather early to make relevant statements, given the limited understanding of this magnetoreceptor.
@@
Figure 6: Magnetic sensitivity $\widetilde{\delta B}=\delta B\sqrt{\tau}$ as a function of $L$, the linear dimension of the sensor, for several mechanisms considered in animal magnetoreception. The red solid line defines the energy resolution limit given by $\widetilde{\delta B}=\sqrt{2\mu_{0}\hbar/L^{3}}$. (a) Cryptochrome magnetoreception based on the FAD-Tr radical-pair sits right at the limit when $2\times 10^{5}$ cryptochromes contribute to a magnetic sensitivity of $\delta B=0.3~{}\upmu{\rm T}$ within time $\tau=10~{}\upmu{\rm s}$. The shaded region accounts for a range of $\tau$ given in terms of the dimensionless parameter $\tilde{\tau}$, i.e. $\tau=10\tilde{\tau}~{}\upmu{\rm s}$. The number of cryptochromes required to hit the ERL limit of $\hbar$ is a function of $\tilde{\tau}$, i.e. $N_{\rm cry}^{0}(\tilde{\tau})=2\times 10^{5}/\tilde{\tau}$. If the number of contributing cryptochromes is e.g. $10^{3}$ times larger than $N_{\rm cry}^{0}(\tilde{\tau})$, the total volume of the radical-pair sensor, and thus the ER, also increase by a factor $10^{3}$, while the linear dimension increases by 10. This is depicted by the green/blue arrows pointing towards increasing linear dimension. The blue arrow shows the ER in units of $\hbar$, and the green arrow the ratio $N_{\rm cry}/N_{\rm cry}^{0}(\tilde{\tau})$. (b) Similar comments apply to the MagR sensor, shown to hit the ERL when $N_{\rm MagR}^{0}(\tilde{\tau})=10/\tilde{\tau}$ MagR-protein complexes contribute to detecting the same $\delta B$ within $\tau=100\tilde{\tau}~{}\upmu{\rm s}$, where again $\tilde{\tau}$ is a dimensionless parameter. For the magnetite mechanism involving a chain of superparamagnetic clusters it is $\widetilde{\delta B}=5~{}{\rm nT/\sqrt{Hz}}$ for a linear dimension of $3~{}\upmu{\rm m}$. (c) The laser spectroscopic measurement of CPF takes place in a solution of 3 mm linear dimension, with the corresponding sensitivity being $\widetilde{\delta B}=2.5~{}{\rm nT/\sqrt{Hz}}$. Picture of cryptochrome protein adapted from [149] under the Creative Commons Attribution License.
@@
Overall, the values for $\widetilde{\delta B}=\delta B\sqrt{\tau}$ and the spatial dimension $L$ that we will use for MagR magnetoreception in our global results of Sec. VII are $\widetilde{\delta B}\approx 3\sqrt{\tilde{\tau}}~{}{\rm nT/\sqrt{Hz}}$ and $L=(N_{\rm MagR}V)^{1/3}\geq 31/\tilde{\tau}^{1/3}~{}{\rm nm}$, where $N_{\rm MagR}$ is the number of MagR complexes contributing to the estimate of the magnetic field, and necessarily, it is $N_{\rm MagR}\geq N_{\rm MagR}^{0}(\tilde{\tau})$.